# CREW: Facilitating Human-AI Teaming Research

### v1

@@ 3.7 Modular Pipeline Design for Quantifying Human Characteristics
Individual differences among humans can significantly affect their teaming with AI agents. To support research along this line, CREW supports a set of cognitive tests to quantify these differences shown in Tab. 4. We provide a modular and convenient pipeline (Fig. 7) for executing cognitive tests and Human-AI experiments. The framework integrates various media files (e.g., instruction videos or pictures), inter-trial intervals, executable Python scripts, and Unity builds, ensuring a smooth and effective workflow. The pipeline requires minimal effort from researchers during proctoring, as a single click initiates the sequential execution of experiments. The pipeline allows restart from interruption points.
@@ 4 Benchmarking Study
As an example of running experiments with CREW, we benchmark a state-of-the-art real-time human-guided RL framework, Deep TAMER [55], along with strong RL baselines. In the original Deep TAMER, the framework was only tested on Atari Bowling with 9 human subjects. With CREW, this is the first time it is possible to systematically conduct human-guided RL benchmarking across multiple environments on a larger population. We summarize our findings as well as insights on the scalability of the framework in this section. We also discuss the relationship between human characteristics and guided agent performance.
@@
Theory of Behavior (theory) [16]
@@
Mental Rotation (rotation) [47]
@@
Mental Fitting (fitness) [47]
@@
Spatial Mapping (spatial) [10]
@@ 4.1 Experiment Setup
Human Trainers We recruited 50 human subjects for the experiments under the approval of the Institutional Review Board. For each 1.5-hour session, the subject will receive \$$\$20$ compensation. Benefiting from CREW’s unique feature to host experiments with parallel sessions, we were able to conduct all 50 experiments within one week. Prior to guiding the agents, the participants were asked to complete all our cognitive tests in Tab. 4.
@@
Deep TAMER leverages human feedback as value functions. During training, a human trainer provides positive or negative discrete feedback based on the agent’s behavior. This feedback is assigned to relevant state-action pairs through a credit assignment window. A neural network is trained to predict human feedback given state-action pairs, and the policy chooses actions that maximize this predicted feedback. Originally, Deep TAMER was limited to discrete actions. We extend it to continuous action spaces using an actor-critic framework, termed as c-Deep TAMER. For bowling, the agent is trained for 5 minutes, and for 10 minutes in Find Treasure and Hide-and-Seek.
Baseline RL We selected two state-of-the-art RL algorithms: Deep Deterministic Policy Gradient (DDPG) [34] and Soft Actor-Critic (SAC) [23].
Heuristic feedback We also evaluate simulated feedback-guided RL. We simply add the feedback signals as additional dense rewards to the environment reward. DDPG is selected as the backbone RL algorithm as the state-of-the-art transitioned from SAC to DDPG [58] in visual control tasks.
4.2 Results
Agent training performance We hypothesize that subjects with higher cognitive tests can lead to higher-performing agents. Therefore, we show the agent performances guided by the 15 subjects who scored the highest in our cognitive tests side by side with the performance of all 50 subjects in Fig. 8. As shown in the results, the agents guided by the top 15 subjects exhibit higher performance than the overall average. In particular for the top 15 subjects, on the simple bowling task, c-Deep TAMER surpasses RL baselines by an average of 10 scores given the same training time. On Find Treasure, heuristic feedback achieved the highest performance as expected, showing the upper bound performance with accurate and non-delayed dense rewards. c-Deep TAMER also demonstrated strong performance with faster learning trends than RL baselines. On 1v1 Hide-and-Seek, c-Deep TAMER performed similarly to RL baselines, suggesting that c-Deep TAMER has difficulty scaling to tasks with higher complexity. Similar conclusions still hold for all 50 subjects.
Figure 8: The average result of the human subjects with the top 15 cognitive test scores and the average result across all 50 subjects. Subjects scoring higher on cognitive tests are generally better at guiding RL. We also find that c-Deep TAMER has difficulties scaling to more complex tasks.
Analysis of Individual Differences Due to the cognitive test feature and emphasis on Human-AI teaming, we can deepen our understanding of how individual human differences can affect the performance of human-guided agents. We calculated the correlation between human subjects’ cognitive test scores and c-Deep TAMER training results in Fig. 9. The cognitive test scores are normalized by the mean and variance over the subjects through z-score, and outliers with scores 1.5$\times$ Interquartile Range above the third quartile or below the first quartile are removed. We found that non-outlier subjects performed equally well on the spatial mapping test so we do not include this for correlation analysis. The heatmap shows the sign of correlation coefficients from linear regression and the statistical significance (*).
Figure 9: Correlation between cognitive test scores and c-Deep TAMER training performance. The darker the color, the more statistically significant the correlation. “+” or “-” means positive or negative correlation.
5 Conclusion, Limitation, and Future Work
6 Acknowledgements
We thank the members at Duke General Robotics Lab, Sida Zhu and Harish Yerra for early development on the platform, and Jiaxun Liu and Easop Lee for beta testing. This work is supported in part by ARL STRONG program under awards W911NF2320182 and W911NF2220113.
Appendix A Accessing CREW
CREW is a platform for cross-disciplinary human-AI teaming research. The codebase for our platform, including environments, interfaces, and algorithms can be accessed through https://github.com/generalroboticslab/CREW. Videos, documentation, and tutorial are hosted on http://generalroboticslab.com/CREW. Our platform is fully open-sourced for academic use.
Appendix B Platform details
The environments of CREW is implemented using Unity 2021.3.24f1, with packages ML Agents 2.3.0-exp.3 [28], Netcode for GameObjects 1.3. [2] and Nakama Unity 3.6.0. [1]. Algorithms are developed with torchrl 0.3.0 [12].
Appendix C Additional Results
C.1 Feedback Visualization
In Fig 10 we show examples of the state-action pairs and human assigned feedback value during c-Deep TAMER training. Since humans trainers do not see the action (next navigation destination) explicitly, we show consecutive frames instead.
Figure 10: c-Deep TAMER training examples.
C.2 Full Cognitive Test Analysis
We include the full results for the cognitive test and c-Deep TAMER score analysis. All linear regression plots including coefficients and p-values and summarized in Fig 11.
Figure 11: Linear Regression Plots of Correlation Analysis
Appendix D Computational Resources
All human subject experiments were conducted on desktops with one NVIDIA RTX 4080 GPU. All evaluations were run on a headless server with 8 $\times$ NVIDIA RTX A6000 and NVIDIA RTX 3090 Ti.
Appendix E Benchmarking Experiment Details
E.1 Implementation details
c-Deep TAMER Algorithm The detailed c-Deep TAMER algorithm is summarized in Alg 1. It is modified from the original Deep TAMER integrated with an actor network for action selection. The actor is updated by gradient ascent that outputs the actions that maximizes the predicted human feedback.
@@ E.1 Implementation details
1:initial policy $A$, feedback model $\hat{H}$, empty replay buffer $\mathcal{D}$, step size $\eta$, buffer update interval $b$, credit assignment window $w$
@@
5:     execute action $a_{i}=\text{clip}(A(\mathbf{s}_{i})+\epsilon,a_{Low},a_{High})$, where $\epsilon\sim\mathcal{N}$
@@
12:         update $\hat{H}$ by one step of gradient descent using
$\nabla\hat{H}\frac{1}{|\mathcal{D}_{j}|}\sum_{(\mathbf{x},\mathbf{y})\in\mathcal{D}_{j}}(\hat{H}(s,a)-y)^{2}$
@@
$\nabla A\frac{1}{|\mathcal{D}_{j}|}\sum_{(\mathbf{x},\mathbf{y})\in\mathcal{D}_{j}}(\hat{H})(s,A(s))$
@@
19:         update $\hat{H}$ by one step of gradient descent using
$\nabla\hat{H}\frac{1}{|\mathcal{D}_{j}|}\sum_{(\mathbf{x},\mathbf{y})\in\mathcal{D}_{j}}(\hat{H}(s,a)-y)^{2}$
@@
$\nabla A\frac{1}{|\mathcal{D}_{j}|}\sum_{(\mathbf{x},\mathbf{y})\in\mathcal{D}_{j}}(\hat{H})(s,A(s))$
@@
Processing Inputs All tasks in our experiments directly learn from pixels. The frames rendered from the environment are first resized to 100$\times$100 pixels. For Find Treasure and 1v1 Hide-and-Seek, we stack the past 3 frames as input. We then apply simple random shift to the frame stack, as it has been shown to be an effective augmentation strategy for visual reinforcement learning. The shifting factors along the height and width are uniformly sampled from [0, 0.08].
@@
Model architecture We use a 3-layer CNN as the vision encoder, each having 64 channels and followed by batch normalization and ReLU activation function. All actor-critic frameworks uses a 3-layer MLP with 256 neurons in each layer for both the actor and critic network.
E.2 Hyperparameters
The hyperparameter settings for out experiments is summarized in Tab 5.
Table 5: Hyperparameters
c-Deep TAMER DDPG SAC
$\gamma$ 0.99 0.99 0.99
learning rate 1e-4 1e-4 1e-4
max_grad_norm 0.1 0.1 0.1
batch size 16 240 240
frames per batch 8 240 240
alpha_init - - 0.1
target entropy - - -6.0
actor scale_lb - - 1e-4
# Q value nets - 2 2
target update polyak 0.995 0.995 0.995
[…]
credit assignment window bowling[0.2, 4], others[0.2, 1] - -
E.3 Human subject experiment details
Overview We recruit the human subject on campus by using flyers and emails. For every human subject, the experiment time is approximately 40 minutes without interruptions. The experiment starts with cognitive tests (10 minutes) and is followed by the human guiding the agent using the c-Deep TAMER framework (30 minutes). The order of the cognitive tests is Eye Alignment, Reflex, Theory of Behavior, Mental Rotation, Mental Fitting, and Spatial Mapping. There are detailed instruction videos for each test before the test starts. As for the human guiding agent part, each human subject guides the agent to play 3 games for a total of 30 minutes(5 minutes for Bowling, 10 minutes for Find Treasure, and 10 minutes for 1v1 Hide-and-Seek). Before each game, there is a detailed instruction video that describes the rule of the game and how human subjects can give feedback to the agent, which we included below.
E.3.1 Instruction Video Script
General Introduction:
Welcome to the General Robotics Lab’s Human AI Collaboration Experiment. Today, our session will start with preliminary cognitive and gaming proficiency tests to gauge your initial skills. Following this, we will delve into the main experiments where you will interact with AI algorithms through a series of engaging games. Our session will conclude with a short survey to capture your feedback on the experience. Your participation is invaluable in advancing our understanding of human AI interactions. Thank you for joining us today, and let’s begin.
Cognitive Test Introduction:
Welcome to the Cognitive Test segment of our experiment. In this session, you will participate in five interactive games designed to assess various cognitive skills for about seven minutes. These tests will challenge your precision, reflexes, predictive abilities, problem-solving skills, and spatial awareness through engaging activities. Each test is brief, and you’ll receive clear instructions before each one begins. Between each trial of each game, there will be a three-second intertrial interval where the computer screen will turn white with a Gray cross in the middle. Please focus on the center of the cross as much as possible during this time. Let’s get started and see how you do.
Eye Alignment Instruction:
In this experiment, your goal is to align the ball positioned on the left side of the screen with the square on the right as accurately as possible. Each trial lasts for five seconds, and you’ll have a total of six trials. Use your mouse to drag the ball during each trial. The time bar at the top center of the screen shows how much time you have for each trail. As time goes down, the bar gets smaller and changes color from green to red.
Reflex Instruction:
In this experiment, after a three-second countdown, the screen turns yellow. You must click as quickly as possible when the screen changes from yellow to green. Clicking during the yellow phase will result in a failure, and failing to click within two seconds after the screen turns green will also fail. You’ll have a total of 6 trials.
Theory of Behavior Instruction:
In this experiment, watch the red ball moving for five seconds, then guess where it will be one second after it pauses. Click on the map to mark your prediction. You have only one chance to click. The closer you are, the higher your score. Remember, you only have two seconds to make your guess after the ball pauses. There will be 6 trials in total. The time bar at the top left shows the time for each trail’s observation and prediction as time decreases. The bar shrinks and changes from green to red.
Mention Rotation and Mental Fitting Instruction:
In this experiment, you will need to answer 12 questions in total, each with only one correct answer. Click the button to choose the option that you think best answers the question. For each question, you have 8 seconds to view and respond. The time bar at the top right of the screen indicates how much time you have for each question. As time decreases, the bar shrinks and changes color from green to red.
Spatial Mapping Instruction:
In this experiment, you will need to answer six questions in total, each with only one correct answer. Watch the video and click the button to choose the option that you think best answers the question. For each question, you are free to answer while the video is playing, and you will also have three extra seconds after the video is paused to answer the question. The time bar at the top right of the screen indicates how much time you have for each question. As time decreases, the bar shrinks and changes color from green to red.
Human Guiding AI Introduction:
Welcome to the Human-Guiding AI section of the experiment. You will guide the AI to play three games for about 30 minutes in total in this section. You’ll give feedback to the AI agents based on their performance in each game. Before starting each game, please enter your name in the name box and click on the Connect button. A match list will appear. Click on the Match button in the list. It will turn brown when clicked. Then click the Join Match button to enter the match. Once you’ve joined the match, select the AI agent in the Active Clients panel at the top right corner of the screen by clicking on it. The agent button will turn brown to indicate that you’ve selected it successfully. Now you can observe the agents and provide feedback. You’ll constantly click on the positive and negative buttons to inform the agent of its performance. Positive means it’s doing well. Negative means it’s doing poorly.
Bowling Instruction:
Each episode of this bowling game consists of 10 rolls. At the start of each roll, there are 10 pins positioned on the right side of the screen. The agent then launches a ball in an attempt to knock down the pins, with the number of pins knocked down determining the score for that roll. The total score for the episode is calculated as the sum of the scores from all 10 rolls. Please guide the AI agent to maximize the total score across all rolls.
Find Treasure Instruction:
In this game, the AI agent, represented by the green character, begins with a limited field of view, only able to perceive surroundings, while the rest of the map is obscured by shadow. As the agent navigates the map, areas it has visited become revealed within its field of view. Please guide the AI agent to locate the treasure and navigate to it, represented by the brown chest, as quickly as possible.
1v1 Hide-and-Seek Instruction:
In this game, the AI agent, represented by the red character, begins with a limited field of view, only able to perceive surroundings, while the rest of the map is obscured by shadow. As the agent navigates the map, areas it has visited become revealed within its field of view. Guide the AI agent to chase the hider, represented by the green character, and catch it as quickly as possible.
E.3.2 Compensation
We pay each human subject \$20 for participation.
Appendix F Author Statement
The authors bear all responsibility in case of violation of rights.

### v3

@@ 3.7 Modular Pipeline Design for Quantifying Human Characteristics
Individual differences among humans can significantly affect their teaming with AI agents. To support research along this line, CREW supports a set of cognitive tests to quantify these differences shown in Table. 3. We provide a modular and convenient pipeline (Figure. 8) for executing cognitive tests and Human-AI experiments. The framework integrates various media files (e.g., instruction videos or pictures), inter-trial intervals, executable Python scripts, and Unity builds, ensuring a smooth and effective workflow. The pipeline requires minimal effort from researchers during proctoring, as a single click initiates the sequential execution of experiments. The pipeline allows restart from interruption points.
@@ 4 Benchmarking Study
As an example of running experiments with CREW, we benchmark a state-of-the-art real-time human-guided RL framework, Deep TAMER Warnell et al. (2018), along with strong RL baselines. In the original Deep TAMER, the framework was only tested on Atari Bowling with 9 human subjects. With CREW, this is the first time it is possible to systematically conduct human-guided RL benchmarking across multiple environments on a larger population. We summarize our findings as well as insights on the scalability of the framework in this section. We also discuss the relationship between human characteristics and guided agent performance.
@@
Theory of Behavior (theory) Chen et al. (2021b)
@@
Mental Rotation (rotation) Shepard & Metzler (1971)
@@
Mental Fitting (fitness) Shepard & Metzler (1971)
@@
Spatial Mapping (spatial) Berkowitz et al. (2021)
@@ 4.1 Experiment Setup
Human Trainers We recruited 50 human subjects for the experiments under the approval of the Institutional Review Board. We highlight that our experiment is the largest-scale study so far on real-time human-guided AI training. For every human subject, the experiment time is approximately 40 minutes without interruptions. The experiment starts with cognitive tests (10 minutes) and is followed by the human guiding the agent using the c-Deep TAMER (which will be discussed in the next section) framework (30 minutes). The order of the cognitive tests is Eye Alignment, Reflex, Theory of Behavior, Mental Rotation, Mental Fitting, and Spatial Mapping. There are detailed instruction videos for each test before the test starts. As for the human guiding agent component, each human subject guides a single agent to play 3 tasks for a total of 30 minutes(5 minutes for Bowling, 10 minutes for Find Treasure, and 10 minutes for 1v1 Hide-and-Seek). Before each task, there is a detailed instruction video that describes the rule of the task and how human subjects can give feedback to the agent. Benefiting from CREW’s unique feature to host experiments with parallel sessions, we were able to conduct all 50 experiments within one week. Prior to guiding the agents, the participants were asked to complete all our cognitive tests in Table. 3.
@@
Deep TAMER Warnell et al. (2018) is a well-established baseline that leverages human feedback as myopic value functions. During training, a human trainer provides positive or negative discrete feedback based on the agent’s behavior. This feedback is assigned to relevant state-action pairs through a credit assignment window. A neural network is trained to predict human feedback given state-action pairs, and the policy chooses actions that maximize this predicted feedback. Originally, Deep TAMER was limited to discrete actions. We extend it to continuous action spaces using an actor-critic framework, termed as c-Deep TAMER, specified in Alg 1. For bowling, the agent is trained for 5 minutes, and for 10 minutes in Find Treasure and Hide-and-Seek.
@@
1:initial policy $A$ parameters $\theta$, human feedback model $H$ parameters $\phi$, empty replay buffer $\mathcal{D}$, step size $\eta$, buffer update interval $b$, credit assignment window $w$
@@
5:    execute action $a_{i}=\text{clip}(A_{\theta}(\mathbf{s}_{i})+\epsilon,a_{Low},a_{High})$, where $\epsilon\sim\mathcal{N}$
@@
12:         update $H$ by one step of gradient descent using
$\nabla_{\phi}\frac{1}{|\mathcal{D}_{j}|}\sum_{(\mathbf{x},\mathbf{y})\in\mathcal{D}_{j}}||H_{\phi}(s_{i},a_{i})-\mathbf{y}||_{2}$
@@
$\nabla_{\theta}\frac{1}{|\mathcal{D}_{j}|}\sum_{(\mathbf{x},\mathbf{y})\in\mathcal{D}_{j}}H_{\phi}(s_{i},A_{\theta}(s_{i}))$
@@
19:         update $H$ by one step of gradient descent using
$\nabla_{\phi}\frac{1}{|B|}\sum_{(\mathbf{x},\mathbf{y})\in B}||H_{\phi}(s_{i},a_{i})-\mathbf{y}||_{2}$
@@
$\nabla_{\theta}\frac{1}{|B|}\sum_{(\mathbf{x},\mathbf{y})\in B}H_{\phi}(s_{i},A_{\theta}(s_{i}))$
@@
Baseline RL We selected two state-of-the-art RL algorithms: Deep Deterministic Policy Gradient (DDPG) Lillicrap et al. (2015) and Soft Actor-Critic (SAC) Haarnoja et al. (2018).
Heuristic feedback We also evaluate simulated feedback-guided RL. We simply add the feedback signals as additional dense rewards to the environment reward. DDPG is selected as the backbone RL algorithm as the state-of-the-art transitioned from SAC to DDPG Yarats et al. (2021) in visual control tasks. The hyperparameters for the heuristic baseline is set to the same as DDPG baseline.
@@
Processing Inputs All tasks in our experiments directly learn from pixels where humans and AI agents share the same input modality. The frames rendered from the environment are first resized to 100$\times$100 pixels. For Find Treasure and 1v1 Hide-and-Seek, we stack the past 3 frames as input to include the history information. We then apply a simple random shift to the frame stack, as it has been shown to be an effective augmentation strategy for visual reinforcement learning. The shifting factors along the height and width are uniformly sampled from [0, 0.08].
@@
Model architecture We use a 3-layer CNN as the vision encoder, each having 64 channels and followed by batch normalization and ReLU activation function. All actor-critic frameworks use a 3-layer MLP with 256 neurons in each layer for both the actor and critic network.
Computational Resources All human subject experiments were conducted on desktops with one NVIDIA RTX 4080 GPU. All evaluations were run on a headless server with 8 $\times$ NVIDIA RTX A6000 and NVIDIA RTX 3090 Ti. We note that we used larger GPU servers due to the need for parallel evaluation resulting from the large scale of our study, but this is not a requirement to run CREW.
4.2 Results
Agent training performance We hypothesize that subjects with higher cognitive tests can lead to higher-performing agents. Therefore, we show the agent performances guided by the 15 subjects who scored the highest in our cognitive tests side by side with the performance of all 50 subjects in Figure. 9. As shown in the results, the agents guided by the top 15 subjects exhibit higher performance than the overall average. This is most prominent in Find Treasure, where c-Deep TAMER (blue curve) is significantly higher in the Top 15 plot than the All 50 plot. As shown in Figure. 10, the correlation between cognitive test scores and guided AI performance is positive for all tasks, where Find Treasure had a strong statistical significance (**, p < 0.01). We do not draw a general conclusion from these early explorations. We hope to provide a useful metric and experiment to study individual differences for future work, exemplified by our benchmarking results. We highlight our concurrent work Zhang et al. (2024) to use them in human-guided RL evaluation.
Figure 9: The average result of the human subjects with the top 15 cognitive test scores and the average result across all 50 subjects. We also find that c-Deep TAMER has difficulties scaling to more complex tasks.
In particular for the top 15 subjects, on the simple bowling task, c-Deep TAMER surpasses RL baselines by an average of 10 scores given the same training time. On Find Treasure, heuristic feedback achieved the highest performance as expected, showing the upper bound performance with accurate and non-delayed dense rewards. c-Deep TAMER also demonstrated strong performance with faster learning trends than RL baselines. On 1v1 Hide-and-Seek, c-Deep TAMER performed similarly to RL baselines, suggesting that c-Deep TAMER has difficulty scaling to tasks with higher complexity. Similar conclusions still hold for all 50 subjects. We also include a comparison between the top and bottom 25 cognitive test scorers in Appendix 7.5.
ginfigure[b!]
Figure 10: Correlation between cognitive test scores and c-Deep TAMER training performance. The darker the color, the more statistically significant the correlation. “+” or “-” means positive or negative correlation. Overall, the total score (i.e., the sum of three game scores) has significant positive correlations with rotation, fitness, and overall score. This is likely a result of the strong contribution of Find Treasure since other tasks only show positive correlations with certain cognitive scores but not statistically significant.
Analysis of Individual Differences Due to the cognitive test feature and emphasis on Human-AI teaming, we can deepen our understanding of how individual human differences can affect the performance of human-guided agents. We calculated the correlation between human subjects’ cognitive test scores and c-Deep TAMER training results in Figure. 10. The cognitive test scores are normalized by the mean and variance over the subjects through z-score. Following Tukey (1977), we remove outliers with scores 1.5$\times$ Interquartile Range above the third quartile or below the first quartile are removed. We found that non-outlier subjects performed equally well on the spatial mapping test so we do not include this for correlation analysis. The heatmap shows the sign of correlation coefficients from linear regression and the statistical significance (*).
5 Conclusion, Limitation, and Future Work
6 Acknowledgment
This work is supported in part by ARL STRONG program under awards W911NF2320182 and W911NF2220113.